# Compass: A Comprehensive Tool for Accurate and Efficient Molecular Docking in Inference and Fine-Tuning

### v1

@@ 2.3.2 Compass Score
The Compass Score is an essential component of DiffDock’s fine-tuning process, designed to incorporate the Binding Affinity, Strain Energy, and Number of Clashes through a nuanced measurement system using LAN-MSE.
@@
  • Binding Affinity Regularizer: Binding affinity is calculated based on the AA-Score from translational ($tr$), rotational ($rot$), and torsional ($tor$) movements and is compared against the ground truth using LAN-MSE:
@@
    where $\hat{y}_{\text{Binding Affinity}}$ represents the predicted Binding Affinity Energy, and $y_{\text{Binding Affinity}}$ denotes the actual value.
  • Strain Energy Regularizer: Strain energy is determined using PoseCheck for the same movements and compared against the ground truth via LAN-MSE:
@@
  • Steric Clash Regularizer: The number of Steric Clashes is also evaluated using PoseCheck and compared to the ground truth:
@@ 2.3.3 Loss Function
To optimize fine-tuning in DiffDock, the overall loss function incorporates the Compass Score total illustrated in Equation 6, which serves as a Non-Gradient-Tracked Loss function penalizer, alongside the primary DiffDock loss, weighted appropriately:
@@
where $\mathcal{L}_{\text{DiffDock}}$ represents the DiffDock Loss Function, which facilitates updates of translation, rotation, and torsional angles; $w_{\text{DiffDock}}$ is the weight assigned to the total loss in DiffDock; $\mathcal{CS}_{\text{Total}}$ represents the total Compass Score as defined in Equation 6; and $w_{\text{Compass Score}}$ is the weight assigned to the total Compass Score, calculated as $1-w_{\text{DiffDock}}$.
@@ 3 Experiments
This section outlines two primary studies conducted during the experimental phase. The first study aimed to demonstrate the noisiness of the trained PDBBind dataset (see Section 3.1), while the second involved fine-tuning predicted positions using the Compass Score as a penalizer (see Section 3.2).
3.1 Pcb Violation Analysis
The experiments utilized the PDBBind dataset, specifically the preprocessed version employed by the state-of-the-art method, DiffDock. Within these experiments, the COMPASS tool was used to assess the results for binding affinity, strain energy, and the number of clashes. It was observed that a significant portion of the preprocessed ligands exhibited issues, which were identified using PoseCheck modules and AA-Score.
@@ 3.2 Fine-Tuning with Compass Score Penalizer
Instead of fine-tuning the entire dataset, a subset of 900 protein-ligand pairs from the DiffDock validation set was initially selected. Due to structural issues in some preprocessed data, as discussed in Section 3.1 (Details in Appendix A.1), the subset was further narrowed down to approximately 700 samples. Given the relatively long calculation time of Compass Score, 80 protein-ligand pairs were randomly chosen for training and 20 for validation.
For the inference test set, similar structural issues led to the selection of about 261 samples from the preprocessed data.
@@ 3.2.1 Experiment Settings
During the fine-tuning phase, weights for the DiffDock loss function (${w}_{\text{DiffDock}}$) were set at 0.99, 0.9, and 0.75, as discussed in Section 2.3.3. Correspondingly, weights for the Compass Score (${w}_{\text{Compass Score}}$) were determined as $1-{w}_{\text{DiffDock}}$.
Additionally, the hyperparameters used in the DiffDock-L model were employed. Various learning rates—0.1, 0.01, 0.001, and 0.0001—were used to optimize performance.
@@ 4 Results
As outlined in Section 3, two primary experiments were conducted. The results of the first experiment, which focused on the analysis of physical violations in the PDBBind dataset, are detailed in Section 4.1. The outcomes of the second experiment, which involved fine-tuning using the Compass Score as a loss function penalizer, are reported in Section 4.2.
Figure 3: Distribution of PCB properties and violations within the PDBBind ground truth dataset. (A): Distribution of binding affinity across the PDBBind dataset. (B): Distribution of the number of steric clashes between protein-ligand pairs in PDBBind. (C): Distribution of strain energy in ligands from the PDBBind dataset.
4.1 Pcb Violation Analysis
As previously mentioned in Section 3.1, we have analyzed the distributions of the PDBBind ground truth data using the COMPASS tool, focusing on binding affinity, the number of steric clashes between protein-ligand pairs, and the strain energy of ligands, as depicted in Figure 3.
Figure 3A illustrates the distribution of binding affinity within the PDBBind dataset, which has a mean of -8.44 kcal/mol and a standard deviation of 3.76 kcal/mol. In Figure 3B, the mean number of steric clashes is reported as 5.85 with a standard deviation of 4.43; notably, only about 5% of the dataset exhibits no steric clashes, indicating minimal physical violations. Lastly, as shown in Figure 3C, the strain energy of ligands displays a mean of 3.11 and a standard deviation of 6.15.
@@ 4.2 Fine-Tuning with Compass Score Penalizer
As detailed in Section 3.2, the DiffDock model was fine-tuned using the Compass Score as a penalizer for the loss function, with the resultant performance enhancements illustrated in Figure 4.
The baseline DiffDock-L model, comprising 30 million parameters, achieved an accuracy of 11.49% for RMSD below 2 Å and 25.05% for RMSD below 5 Å in selected 261 test set(Appendix A.1). Upon fine-tuning the DiffDock-L model, we observed an increase in accuracy to 11.91% for RMSD below 2 Å and 25.71% for RMSD below 5 Å. Application of our method further enhanced the accuracy to 13.02% for RMSD below 2 Å and 27.16% for RMSD below 5 Å.
Comparatively, standard fine-tuning improved performance by 3.67% under RMSD below 2 Å, whereas our method yielded a 13.3% improvement, representing nearly a 3.67-times increase. For RMSD below 5 Å, standard fine-tuning resulted in a 2.60% performance gain, while our method led to an 8.41% increase, marking approximately a 3.23-times enhancement in fine-tuning efficacy.
Figure 4: Benchmarking DiffDock models with original pretrained (DiffDock-L), fine-tuning (DiffDock-L+FineTuning), and Fine-tuning with Compass Score (DiffDock-L+FineTuning+CompassScore(ours)) in RMSD $<$ 2 Å and RMSD $<$ 5 Å
@@ 5 Discussion
This study conducted two key experiments to evaluate the effectiveness of the Compass Score within the molecular docking landscape using the PDBBind dataset. The first experiment (Section 4.1) analyzed PCB violations, uncovering significant variations in binding affinity, steric clashes, and strain energy, which highlight the inherent complexities and inconsistencies of protein-ligand interactions. The results indicated that a majority of the dataset exhibited steric clashes, with only a minor portion free from such violations, suggesting that even well-curated datasets like PDBBind contain inherent inaccuracies that could affect docking prediction outcomes. This underscores the need for robust modeling techniques that can adapt to the diverse physicochemical & biochemical properties observed in various ligand-protein pairs, emphasizing the challenges in predicting precise docking conformations.
The subsequent application of the Compass Score as a penalized (Section 4.2) significantly improved the accuracy of docking predictions in RMSD metric, demonstrating the effectiveness of incorporating comprehensive scoring functions into the fine-tuning process of molecular docking algorithms. The enhancement in performance metrics, particularly under RMSD thresholds of 2 Å and 5 Å, validates the hypothesis that a nuanced scoring system can refine the precision of docking. This method markedly outperformed standard fine-tuning approaches, nearly tripling the enhancement rates and confirming the value of adaptive adjustments that respond to unique dataset characteristics.
Furthermore, as depicted in Figure  5, RMSDs greater than 5 Å can still result in viable molecular conformations. While lower RMSDs generally correlate with improvements in Binding Affinity, Strain Energy, or Steric Clashes, it is not accurate to assert that a specific RMSD threshold consistently yields better molecular conformations during docking processes. This point is further exemplified by the RMSD discrepancies shown in Figure 3, where despite an RMSD value of 4.75 for the 2si0 PDB ID ligand in the initial fine-tuning epoch, significant violations in PCB are evident, highlighting the RMSD’s limitations.
@@ 6 Conclusion
These findings advocate for the integration of detailed, PCB-informed scores within docking algorithms to elevate predictive accuracy and accurate assessment. Future studies should explore the scalability of such refined scoring mechanisms across broader datasets and consider integrating more complex interaction dynamics to further advance the field of DL-based molecular docking. This approach could potentially lead to breakthroughs in the predictive capabilities of computational models, paving the way for more accurate drug discovery and biomolecular research.
Figure 5: Comparison of PDB ID: 5c28 Ligand Comformation predictions during fine-tuning. Ground Truth Binding Affinity, Strain Energy and Number of Clashes indicated as green in the first column
@@ Acknowledgments
AS expresses gratitude to Bora Uyar and Mohammed Maqsood Shaik for their valuable discussions on inference mode fine-tuning mode with Compass.
@@ B.1 AA-Score
AA-score  (Pan et al., 2022) calculates the binding affinity energy as illustrated in Equation 1. Table 1 summarizes energy components like Hydrogen Bond, van der Waals interaction, Electrostatic interaction, Hydrophobic contacts, $\pi-\pi$ interaction, $\pi$-cation interaction and Metal-Ligand Interactions.
@@
Binding Affinity -11.33 3505.32
Strain Energy 7.31 20.65
Number of Clashes 6 205

### v2

@@ 2.4.3 Compass Score
The Compass Score is a key element of the Fine-Tuning mode in CompassDock. It compares the PCB features (binding affinity, strain energy, and number of steric clashes) of the sampled molecular conformation with those of the ground truth conformation using the LAN-MSE loss function.
@@
  • Binding Affinity Regularizer: Binding affinity is calculated based on the translational ($tr$), rotational ($rot$), and torsional ($tor$) movements and is compared to the ground truth using the LAN-MSE loss function:
@@
    where $\hat{y}_{\text{Binding Affinity}}$ represents the predicted binding affinity energy, and $y_{\text{Binding Affinity}}$ denotes the true binding affinity value.
  • Strain Energy Regularizer: Strain energy is calculated based on the same movements and compared to the ground truth strain energy using the LAN-MSE loss function:
@@
  • Steric Clash Regularizer: The number of steric clashes is calculated based on the same movements and compared to the ground truth using LAN-MSE:
@@ 2.4.4 Loss Function
To optimize CompassDock in Fine-Tuning mode, the overall loss function incorporates the total Compass Score, shown in Equation 6, which acts as a non-gradient-tracked penalizer alongside the primary DiffDock loss, with both components weighted accordingly:
@@
where $\mathcal{L}_{\text{DiffDock}}$ represents the DiffDock loss function, which updates translation, rotation, and torsional angles; $w_{\text{DiffDock}}$ is the weight assigned to the DiffDock loss; $\mathcal{CS}_{\text{Total}}$ represents the total Compass Score (as defined in Equation 6); $w_{\text{Compass Score}}$ is the weight assigned to the total Compass Score, calculated as $1-w_{\text{DiffDock}}$.
@@ 3 Experiments
3.1 PCB Quality Analysis
The experiments used the PDBBind dataset, specifically the preprocessed version applied by DiffDock. In experiments, the unified Compass workflow was employed to evaluate binding affinity, strain energy, and the number of clashes. A notable portion of the preprocessed ligands were found to have unfavorable PCB features.
@@ 3.2 Fine-Tuning with Compass Score Penalizer
Instead of fine-tuning the entire dataset, an initial subset of 900 protein-ligand pairs from the PDBBind validation set was selected. Due to structural issues in some of the preprocessed data, as discussed in Section 3.1 (Appendix A.1), this subset was further reduced to approximately 700 samples. Given the relatively long computation time for the Compass Score, 80 protein-ligand pairs were randomly selected for training and 20 for validation.
For the test set, similar structural issues resulted in the selection of about 261 samples from the preprocessed data.
@@ 3.2.1 Experiment Settings
During the fine-tuning, the weights for the DiffDock loss function (${w}_{\text{DiffDock}}$) were set at 0.99, 0.9, and 0.75 (Section 2.4.4). The corresponding weights for the Compass Score (${w}_{\text{Compass Score}}$) were determined as $1-{w}_{\text{DiffDock}}$.
Additionally, the hyperparameters from the DiffDock-L model were applied, and various learning rates—0.1, 0.01, 0.001, and 0.0001—were used to optimize performance.
@@ 4 Results
Figure 3: Distribution of PCB properties and quality within the PDBBind ground truth dataset. (A): Distribution of binding affinity across the PDBBind dataset. (B): Distribution of the number of steric clashes between protein-ligand pairs in PDBBind. (C): Distribution of strain energy in ligands from the PDBBind dataset.
4.1 PCB Quality Analysis
We analyzed the PDBBind ground truth data using the unified Compass workflow (Section 3.1), focusing on binding affinity, steric clashes between protein-ligand pairs, and ligand strain energy, as illustrated in Figure 3.
Figure 3A shows the distribution of binding affinity in the PDBBind dataset, with a mean of -8.44 kcal/mol and a standard deviation of 3.76 kcal/mol. In Figure 3B, the average number of steric clashes is 5.85, with a standard deviation of 4.43; notably, only about 5% of the dataset has no steric clashes, indicating minimal physical violations. Finally, Figure 3C illustrates the distribution of ligand strain energy, with a mean of 3.11 and a standard deviation of 6.15.
@@ 4.2 Fine-Tuning with Compass Score Penalizer
Figure 4: Benchmarking DiffDock models: pre-trained model (DiffDock-L), fine-tuned model (DiffDock-L+FineTuning), and fine-tuning with Compass Score (CompassDock (ours)), based on RMSD and favorable PCB properties.
The performance of CompassDock in Fine-Tuning mode was benchmarked (Section 3.2) against DiffDock-L in inference and DiffDock with fine-tuning based on the percentage of ligands with RMSD $<$ 2Å and RMSD $<$ 5Å, as well as the percentage of ligands with favorable PCB features (binding affinity $<$ 0, strain energy $<$ 5, and number of steric clashes $<$ 5), as shown in Figure 4.
The baseline DiffDock-L model, consisting of 30 million parameters, achieved an accuracy of 11.38% for RMSD below 2 Å and 24.67% for RMSD below 5 Å on the selected 261-sample test set (Appendix A.1). After fine-tuning, the DiffDock-L model improved slightly, reaching 11.57% accuracy for RMSD below 2 Å and 25.75% for RMSD below 5 Å. Applying the Compass Score method yielded a modest enhancement, with 11.42% accuracy for RMSD below 2 Å and 24.75% for RMSD below 5 Å.
In terms of favorable PCB ligands (defined as binding affinity $<$ 0, strain energy $<$ 5, and number of steric clashes $<$ 5), the original PDBBind test set achieved 41.38%. The DiffDock-L model performed at 4.21%, and fine-tuning further reduced this to 3.83%. However, applying the Compass Score in CompassDock improved the percentage of favorable PCB ligand conformations to 4.98%.
@@ 5 Discussion
This study conducted two key experiments to assess the effectiveness of the Compass Score within the molecular docking framework using the PDBBind dataset. The first experiment (Section 4.1) evaluated PCB feature quality, revealing notable variations in binding affinity, steric clashes, and strain energy, which highlight the complexities and inconsistencies of protein-ligand interactions. The findings showed that most of the dataset exhibited steric clashes, with only a small portion free from such violations. This suggests that even widely-used datasets like PDBBind contain inherent inaccuracies that could influence docking prediction outcomes. These results emphasize the need for robust modeling techniques that can handle the diverse physicochemical and biochemical properties found in various ligand-protein pairs, underscoring the challenges in predicting accurate docking conformations. This aligns with the approach taken by the PLINDER dataset  (durairaj2024plinder).
Figure 5: Comparison of PDB ID: 5c28 Ligand Conformation predictions during fine-tuning. Green Structures: Ground Truth Conformation of Ligand; Cyan Structures: Ground Truth Structures of Binding Pocket. Ground Truth Binding Affinity, Strain Energy, and Number of Clashes indicated as green in the first column
Although the use of the Compass Score as a loss penalizer (Section 4.2) resulted in only a slight improvement in RMSD accuracy compared to standard fine-tuning, our results showed that standard fine-tuning tends to reduce the capture of favorable PCB properties. In contrast, incorporating the Compass Score improved the favorable PCB characteristics. This highlights the effectiveness of integrating PCB properties into the fine-tuning process of molecular docking algorithms.
Furthermore, as shown in Figure  5, RMSDs greater than 5 Å can still result in viable molecular conformations. While lower RMSDs correlate  (Alhossary et al., 2015; Hassan et al., 2017; McNutt et al., 2021) with improvements in binding affinity, strain energy, or steric clashes, it is not accurate to assert that a specific RMSD threshold consistently leads to better molecular conformations during docking processes. This point is further exemplified by the RMSD discrepancies shown in Figure 3, where despite an RMSD value of 4.75 for the 2si0 PDB ID ligand in the initial fine-tuning epoch, significant violations in PCB are evident, highlighting the RMSD’s limitations.
Binding Affinity -11.33 3505.32
Strain Energy 7.31 20.65
Number of Clashes 6 205
@@ 6 Conclusion
These findings support the integration of detailed PCB features into docking algorithms to improve predictive accuracy and assessment. Future research should investigate the scalability of refined scoring mechanisms across larger datasets and explore the incorporation of more complex interaction dynamics to further advance DL-based molecular docking. This approach holds the potential to significantly enhance the predictive capabilities of computational models, leading to more accurate outcomes in drug discovery and biomolecular research.
@@ Acknowledgments
AS expresses gratitude to Mohammed Maqsood Shaik for their valuable discussions. We also thank Dan Munteanu, Ricardo Wurmus and Martin Siegert for their help with the compute servers.
@@ B.1 AA-Score
AA-score  (Pan et al., 2022) calculates the binding affinity energy as illustrated in Equation 8:
$\sum_{i}\Delta G_{\text{vdW--main}}^{AA}$
Table 1 summarizes energy components like Hydrogen Bond, van der Waals interaction, Electrostatic interaction, Hydrophobic contacts, $\pi-\pi$ interaction, $\pi$-cation interaction and Metal-Ligand Interactions.